Grade 12th passMechanics

A ship of mass 3×10^7 kg is at rest.By applying a force of 30×10^5 Newton it is displaced through a distance 25 m. Then the speed of ship after this displacement is

Supriya
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student
 
 a = F/m = (30×105)/(3×107 ) = 0.1 m /sec^2
Speed attained by the ship, v = √2as = 2.236 m/s
 
 
Regards
Arun (askIITians forum expert)